Would you please help me to solve this:
For the line with the equation below:
a. Transform the equation into slope-intercept form.
b. Find the slope and y-intercept of the line.
c. Find the equation, in point-slope form, of the line that is perpendicular to this line
and passes through the point (1, 2).
4x – 8y = 8
Thanks
(This is the 2nd time that I am submitting this question, can somebody help pls?)

Answer by lwsshak3(11628)

Would you please help me to solve this:
For the line with the equation below:
a.Transform the equation into slope-intercept form.
4x–8y=8
8y=4x-8
y=x/2-1
..
b.Find the slope and y-intercept of the line.
Standard form of equation of a line:y=mx+b, m=slope, b=y-intercept
For given equation:4x–8y=8
slope=1/2
y-intercept=-1
..
c.Find the equation, in point-slope form, of the line that is perpendicular to this line
and passes through the point (1, 2).
slope of perpendicular line=-2 (negative reciprocal of slope of given line)
point-slope form of equation: y1-y2=m(x1-x2)
using coordinates of given point (1,2),
Equation: y-2=-2(x-1)